5. Given C 5000.75 (Y T) I = 2,000 -50r G 1,000 T 1,000 - 10Y - 2000r Ms 50,000 P a. Derive the IS curve and the LM curve, anPlease give a detailed solution, thank you!
Answer #1

(A) Y=C+I+G

Y= 500+0.75(Y-1000) + 2000-50r + 1000

0.25Y= 2750 - 50r (IS)

Md/P = Ms/P

10Y-2000r = 50000 (LM)

EQUATING IS AND LM, WE GET :

Y = 11000-200r

Y= 5000+200r

Y= 8000

r= 15%

(B) Now govt spending is 1500

therefore,

Y= 500+0.75(Y-1000) + 2000-50r + 1500

Y= 13000-200r (is)

Y= 5000+200r (lm)

Y= 9000

r= 20%

if they want to keep the interest rate fixed increase money supply, shift lm curve to the right

y= 5000+ 200r

put 15% r value and you get change in Y= 5030, Y= 13030

similarily if you want to keep the output fixed

earlier Y= 8000

8000-5000= 200r

CHANGE IN r= 15%

r= 15+15= 30%